bulten - 12:43 Wednesday 07 April 2021 (51340)
Print this report
adjusted setpoint SNEB SBE Ty top filter
the bench TY errors were rather large; for the top filter the control on Ty was at -165 mu rad. I adjusted the Ty setpoint of the SBE top filter and that led to better TY control of the bench.
When closing the SBE loop, suddenly F1-Y LVDT signal jumped from +1.5 mm to 12 micrometer; maybe there is a short in the F1-y LVDT; since it is not part of the control I will leave it like this and try to inspect when the beam is not used (by adjusting the microsprings and seeing if we can trade off F1 and F0 y position). It might also be that the controls really changed the equilibrium point of F1 very abruptly, but the transition was REALLY sharp (see plot). The small excursions in x and z show the settling at the end of closing the loop.
Non-image files attached to this report